OpenStudy (anonymous):

find formula for the inverse of function f(x)=sqrt{10-3x}

OpenStudy (anonymous):

solve \[x=\sqrt{10-3y}\] for \(y\) via \[x^2=10-3y\] \[x^2-10=-3y\] \[y=\frac{x^2-10}{-3}=\frac{10-x^2}{3}\]